# Towards Interpretable Sequence Continuation: Analyzing Shared Circuits in Large Language Models

### v5

@@ 7 Ablating Sequence Continuation Components on Math-Related Prompts
We examine how ablating sequence continuation components affects the generated output for Llama-2-7B on various prompts. This demonstrates how these circuits are not just specific to a narrow range of sequence continuation tasks, but have effects on a wide range of mathematical reasoning tasks.
We ablate important components (ie. reduce performance greatly when ablated). Then, we observe the ablation effects on the following prompt types: (1) Additive sequences for intervals >1 such as "2 4 6", (2) Arithmetic, and (3) Natural language math reasoning prompts. In Table 2, we provide a summary of these experimental results. More results, including those for Spanish sequences and for GPT-2, are in Appendix A.
@@
We measure how well the model correctly performs completions on several prompt types after ablating component sets of our discovered sequence continuation circuits. Given a set of prompts, we define the Percentage Destroyed after Ablation as the percentage of prompts which do not complete correctly after ablation. We compare our component set ablations with random component ablations which have no overlap with the intersection of all three component sets. We use the same 50 randomly chosen component sets for every prompt, and take their mean score.
@@
On intervaled sequences. As shown in Table 2, for intervaled sequences of +2, +3, +10, and +100, we discover that using the entire component set for a task destroys the ability for 96-100% of cases. In contrast, using random component ablations only destroys around 3-28% of cases.
Notably, we are using circuits discovered for +1 interval continuation, we found that these circuits have little overlap with the component sets for +2 and +3 circuits found by the iterative ablation methods. This suggests that there is overlapping functionality between the intervaled sequence continuation circuits.
On arithmetic. Ablating sequence continuation head sets destroy the model’s ability to perform addition and subtraction destroys its ability to perform, but not its ability to perform multiplication. This suggests that sequence continuation, addition, and subtraction tasks share circuits.
On math-related word problems. We ablate the model on math reasoning word problem templates, such as on the template "If this month is X, and Y months pass, what month is it? ". We found that for simpler prompts such as "What are the months in a year? ", we need to ablate all the heads of a sequence continuation circuit for there to be destruction effects that are more notable than several runs of ablating the same number of heads at random. However, for more complex prompts that may seemingly be dependent on almost all the heads of the network, we only need to ablate the 3 very functionally important heads- 5.25, 16.0, and 20.17- for there to be an effect that is more noticeable than ablating 3 heads at random. Appendix A delves into these studies in more detail.
8 Conclusion
9 Limitations
Dataset Size. As there are only twelve months, the number of possible continuing sequences was limited. Additionally, even if months were not used, GPT-2 also has limited prediction ability for number word sequences, as detailed in Appendix F. However, our dataset size for the months continuation task fully captures all of the months; in contrast, a small dataset size brings more issues when it does not capture all of the true distribution.
Future Work. In the future, we plan to dive deeper into this study, such as by performing neuron-level and feature-level analysis, and by examining components exclusive to the number words task to see if they handle mapping between abstract representations of numbers and number words. Our future work plans also include analyzing the effects of model editing on shared, entangled circuits. These may include quantifying the relationship between circuit entanglement and editing impact (which may be done via embedding space projection (Dar et al., 2022)), modifying the sub-circuit used for sub-task $S$ and observing if the ability to recognize $S$ in multiple tasks is destroyed, and utilizing methods such as model steering to edit task $S$ to perform a similar task $S^{\prime}$ Turner et al. (2023); Merullo et al. (2023).
Acknowledgements
We are grateful to Torr Vision Group (TVG) and Apart Lab members for feedback on the previous version, specially, Jishnu Mukhoti, Luke Marks, Clement Neo, Michelle Lo and Ashkan Khakzar. Clement Neo provided the visuals for Figure 1.
Appendix A Ablation on Math-Related Prompts
Zero ablation is performed for these experiments.
Llama-2-7B- Experimental Setup. Unlike in GPT-2 Small, for Llama-2-7B we find most of the MLPs to be necessary to be able to perform sequence continuation. Thus, we do not remove them. As we do not remove MLPs, these results suggest that there are attention heads specific to sequence continuation and arithmetic that are crucial for these operations across various languages. Thus, these results suggest that we do not need a large dataset to obtain the important components via a coarse ACDC variant algorithm.
For random ablation on intervaled sequences and arithmetic, a sequence continuation component set for a task is around 80-90 heads, so we choose 100 heads to ablate, ensuring that they do not include heads from the intersected set of 16 heads of the 3 main tasks.
Overall, we note that these tasks are not solely dependent on the sequence continuation component sets we found, which means random ablations may choose components that are crucial for the task. For instance, the random ablation score for the arithmetic may likely be lower, as some of its runs may be ablating heads that are involved in arithmetic.
@@ Appendix A Ablation on Math-Related Prompts
Llama-2-7B Arithmetic- Extended Analysis. For addition and subtraction tasks, we find that ablating the sequence continuation component heads destroy the ability to perform double-digit arithmetic for 92-100% of our experiments, while random ablated heads only do so  20-33% of the time.
Llama-2-7B Math Word Problems. The following experiments provide some evidence that the sequence continuation circuits are not specific to just sequence continuation, but can extend to prompts involving sequence continuation reasoning. More experiments on larger datasets and more prompts can determine if statement holds in more cases. In summary, we evaluate on the following prompt templates:
  1. 1.
    "What are the months in a year? Give all of them as a list. Be concise."
  2. 2.
    "If this month is X, and Y months pass, what month is it? Answer: "
  3. 3.
    "If today is the Xth of (month M), what date will it be in Y days? "
(1) "What are the months in a year? " We generate 50 tokens from the prompt, "What are the months in a year? Give all of them as a list. Be concise."
When ablating any of the 3 sequence continuation component sets, which contain around 80 heads each (eg. the Numerals circuit has 86 heads), the model loses the ability to correctly complete the answer. In particular, ablating the Numerals head set will make the model stop the list at June, ablating the Number words head set will make the model stop the list at February, and ablating the Months head set will make the model state the output as "1. January 2. February 2. February 3. March 3. April 3. May 3. August 3. August", listing each item on a new line. We note that the Numerals and Number Words head sets do not particularly show strong evidence for sequence destruction, but the Months head set clearly shows the model loses the ability to correctly complete the sequence.
In contrast, choosing 86 random heads that do not overlap with the Numerals circuit component set (which has 86 heads) will allow the model to retain the correct ordering for 47 out of 50 random samples of 86 random heads. We evaluate these outputs using GPT-4.
(2) "If this month is X, and Y months pass, what month is it? " We study ablation on prompts of the format, "If this month is X, and Y months pass, what month is it? ". We found that giving an instruction such as "Be concise" or a pattern we expect the model to complete in, such as "If this month is March, and 2 months pass, what month is it? Answer: May. ", often makes the model give wrong answers. Instead, we found that just letting the model generate from the question itself would beget the right answer many times.
Thus, as the prompt answers are in different formats for each prompt (eg. "12 months have passed, so it is now December again" vs "If this month is April and 1 month passes, then it is May", etc.), we first manually inspect ablation effects on a small set of prompts in which the correct answer is given for the unablated model. We found that using GPT-4 to evaluate often yielded several errors, so its accuracy is not guaranteed to be sound. Thus, when evaluating a larger set of prompts, we combine analysis from a simple string parsing function, manual inspection, and GPT-4 to increase scoring accuracy.
We use 10 randomly sampled component sets, and generate 15 tokens in the model output. We found that for most prompts, this is enough for the model to correctly obtain the right answer, as shown for a sample in Table 3.
We use prompts that the model completes correctly, and discover that we only need to ablate the 3 functionally important heads for sequence continuation- 5.25, 16.0, and 20.17- to disrupt these word prompt correct completion. In contrast, ablating just 4 heads at random often does not destroy the model’s correct completion abilities.
In detail, using just the 3 functionally important heads yields a percentage destroyed score of 42.9%, while using random heads yields a mean destruction score of 7.14% of prompts, for 10 randomly sampled runs of 7 prompts. These heads should work together; if one of them is not present, the model will still retain its correct completion ability.
We also evaluate on 43 other prompts using this template, with 10 runs for random ablations (meaning 430 runs in total), scoring them using GPT-4 (noting that it may not be as accurate as manual inspection; to improve its ability, we provide it with a list of the correct answers along with the generated outputs, and manually inspect to see if it obtains which indices have correct outputs). Additionally, we also use a simple method that just checks if the correct answer appears in the output; however, this may not be as accurate because that month may just be stated, but not stated as the answer.
We find that ablating the 3 functionally important heads yields a destruction score of  37.21%, and the random heads yields a mean destruction score of  14.19%. For each run, we calculate the percentage destroyed score, and take the mean over every run. These scores may not be as accurate however, and may change after revision via further manual inspection; thus, we only use them as a rough estimate.
In contrast, we find that a "simpler" sequence continuation task such as "1 2 3" requires more heads- around 40 to 80- for ability destruction to occur for most prompts. This may suggest that more complex prompts are more sensitive to requiring most components be healthy, while certain simpler tasks do not require this and are only destroyed when many components are destroying, possibly suggesting that these large circuit component sets we discovered may contain many backup sub-circuits within them.
(3) "If today is the Xth of (month M), what date will it be in Y days? " Similar to the previous prompt type, we have the model generate 5 tokens in sequence. However, we do not use GPT-4 to evaluate this correct answer, as the correct answer for the majority of prompts is given in "Month Day" format (eg. January 10th), allowing it to be evaluated algorithmically without advanced large language models.
We use 41 prompts that the model completes correctly, and discover that we only need to ablate the 3 functionally important heads for sequence continuation- 5.25, 16.0, and 20.17- to disrupt these word prompt correct completion. In contrast, ablating just 3 heads at random often does not destroy the model’s correct completion abilities. In detail, we find that ablating our three important heads destroys the correct completions for 24.4% of prompts, while randomly ablating on 3 heads, on average over 10 runs, destroys 9.2% of prompts. We also notice that ablating the 3 functionally important heads often shifts the day by just a few days, rather than changing the month.
We note that the last two prompt types are still sensitive to ablating any 3 heads of the model, suggesting that the model may require most of the heads of the model to work properly for these tasks, or that its predictions for the right answer are not as strong (having more weight on the right answer over other tokens). The latter is likely true given that the model does not always obtain the correct answer for these prompt templates; given 100 prompts, we selected 43 and 41 prompts, respectively, for the previous two templates as these were the ones the model correctly completed on. The destruction score for the 3 heads is also closer to the random ablation score. Thus, this prompt type is relatively weaker evidence for the effect of the discovered sequence continuation circuits on math-reasoning word problems; nonetheless, there is some noticeable effect, so we have reported its results in this Appendix. Using a better (fine-tuned or larger) model that has improved reasoning skills can mitigate this issue.
@@
Single-Digit Addition 62% 46% 92% 15.6%
Single-Digit Subtraction 72% 46% 74% 12.4%
Spanish sequence continuation. We looked into Llama-2-7B’s capabilities for sequence continuation across various languages, but found it to be limited. For instance, the model is able to complete "uno dos tres" correctly, but will output numeral "6" for prompt "dos tres cuatro cinco" instead of "seis". Thus, we only study two prompts for Spanish number word continuation that Llama-2-7B successfully completes on: "uno dos tres" and "seis siete ocho", and two prompts for Spanish months continuation: "enero, febrero, marzo". Lastly, we find the model can also correctly complete Spanish months continuation prompts when asked in natural language question formats, and find that ablating the sequence continuation circuits also affects these inputs.
GPT-2 Small- Extended Analysis. We ablate the sub-circuit, successor head 9.1, and random component(s), and observe the effects on sequence continuation prompts. As GPT-2 Small does not succeed on many math-related prompts, we only tested prompts such as math reasoning in the form of word problems on Llama-2-7B.
While Gould et al. (2023) show that the successor head is crucial for incrementing, our experiments demonstrate that ablating just the successor head is not enough to change the logits and the correct answer for continuing sequences. Instead, ablating the 3 heads and MLP 9 is enough to change the correct answer, and the entire generated sequence afterwards. We compare this to ablating 3 random heads with MLP 9 that do not pass the threshold; these random heads, on average, do not corrupt the model enough.
We also show that most subsets of this combination of 4 components is not enough by themselves. Just ablating MLP 9, for instance, still allows for generation. In many cases, the model requires a combination of this sub-circuit to work; if all of them are ablated, it will fail. But just ablating the successor head doesn’t destroy the model’s ability to work. We hypothesize that there are distributed backups.
Appendix B GPT-2 Small Connectivity Details
As seen in both Figure 9 and in Table 6 in Appendix I, in which only head 5.0 of the Numerals circuit is not part of the Number Words circuit, the Numerals circuit is nearly a subset of the Number Words circuit. This suggests that the Number Words circuit uses the Numerals circuit as a sub-circuit, but requires additional components to make accurate predictions.
Numerals Task NumWords Task Months Task
Numerals Circuit 81.01% 48.41% 113.52%
Number Words Circuit 87.35% 81.11% 103.64%
Months Circuit 43.74% 32.36% 80.30%
IOI Circuit -6.70% -15.82% -9.20%
Appendix C GPT-2 Small Functionality Details
C.1 Last Token Sequence Detection Head
C.2 Successor Head Scores
To check that head 9.1 outputs next sequence tokens, we study its component output scores. Table 5 shows that given a numeral token I as input (eg. 1), head 9.1 often outputs a token I+1 or higher (eg. 2). For numerals between 1 and 100, its next score is 87%, while its copy score is 59%. We also note that the next sequence scores of most heads are low, with an average of 3.29%, and that head 9.1 has the highest next sequence score. Thus, it seems to function as a "next sequence head".
This is reinforced by the next sequence score for number words, which is 90.63%, while the average for all attention heads is 2.97%. Although head 9.1 does not appear to output months given any month token, we observe something peculiar: 9.1 is the only head that will output the next rank given a month (eg. given "February", output "third", and its "next rank given month" score is 31.25%. This appears to be related to how months can be mapped onto ranks.
Token Top-3 Tokens after Unembed
‘78’ ’ 79’, ‘80’, ‘81’
‘six’ ’ seventh’, ’ eighth’, ’ seven’
‘August’ ‘ighth’, ‘eighth’, ‘ninth’
Appendix D Llama-2-7B Functionality Details
As the full circuits are too large to include in this paper, we only discuss important components.
Sequence Continuation: Other Attention Patterns Figures 7 and 8 show attention patterns for heads 11.8 and 16.0, which were found to be common in all three sequence continuation tasks. These have a slight resemblence to the "last token sequence detection head" 7.11 in GPT-2 Small, though the resemblence is not as prominent as the mapping between 4.4 and 5.25, and 9.1 and 20.17, in GPT-2 Small vs Llama-2-7B. Despite head 16.0 being more important than head 11.8, its attention pattern is not as strong.
Figure 6: Llama-2-7B attention pattern for Head 20.17, bearing resemblence to the attention pattern of GPT-2 Small’s Head 9.1 in Figure 3, showing the component pays strong attention to only the most recent sequence member.
Figure 7: Llama-2-7B attention pattern for Head 11.8. At the last token, there appears to be a resemblence to head 7.11, where more recent sequence members than earlier sequence members.
Figure 8: Llama-2-7B attention pattern for Head 16.0. Tokens appears to attend to the last period of the first instance of the pattern (<object> lost in Month.) However, it is not clear from this plot what the role of this head is. Despite being more important than head 11.8, its attention pattern is not as strong.
Appendix E Computational Resources and Packages
@@
For each task, the node and edge iterative methods took a total of 1 to 2 hours to run on an A100 GPU. The code for the experiments was written in Python, utilizing the TransformerLens package, and were run on Colab Pro+.
Appendix F GPT-2 Experimental Setup Details
@@
Appendix G Llama-2 Experimental Setup Details
@@
Dataset. For Llama-2, we ran iterative node pruning on samples consisting only of sequence members, and ran attention patterns on samples consisting of both sequence and non-sequence members. This is because for Llama-2, the main focus was on finding similar components between the sequence continuation sub-circuits of GPT-2 and Llama-2, rather than finding more minimal circuits, which would be beyond the computational resources we allocated.
@@
Iterative Methods. Given that Llama-2 is larger than GPT-2, we modified the iterative component ablation method to first ablate all the heads of an attention layer to see if this makes the model performance perform below the threshold; if it does, we remove all the heads of that attention layer from the candidate circuit. While this method variation is faster if many attention layers are unnecessary, it has the downside in which removing certain unimportant heads may increase performance, which may, in summation, offset the performance degrading abilities of important heads. This means that there may be some important heads which may be missed by this method. However, we found that the most important attention heads were not missed by this method, and the end results were often comparable to the original version of our iterative methods.
Appendix H All Three Circuits Diagram for GPT-2 Small

### v6

@@ 7 Ablating Sequence Continuation Components on Math-Related Prompts
We examine how ablating sequence continuation components affects the generated output for Llama-2-7B on various math-related prompts. This demonstrates how these circuits are not just specific to a narrow range of sequence continuation tasks, but have effects on a wide range of mathematical reasoning tasks. We observe the ablation effects on the following prompt types: (1) Additive sequences for intervals >1 such as "2 4 6", and (2) Arithmetic. In Table 3, we provide a summary of these results. More results, including those for natural language math reasoning prompts and Spanish language sequences, are in Appendix I.
@@
We measure how well the model correctly performs completions on several prompt types after ablating component sets of our discovered sequence continuation circuits. Given a set of prompts, we define the Percentage Destroyed after Ablation as the percentage of prompts which do not complete correctly after ablation. We compare our component set ablations with random component ablations which have no overlap with the intersection of all three component sets. We use 50 randomly chosen component sets for every prompt, and take their mean score. Appendix I discusses details of this ablation procedure.
@@
On intervaled sequences. As shown in Table 3, for intervaled sequences of +2, +3, +10, and +100, we discover that ablating an attention head component set for each sequence continuation task destroys the model’s completion ability for 96-100% of cases. In contrast, using random component ablations only destroys around 3-28% of cases.
@@
On arithmetic. Ablating sequence continuation head sets destroy the model’s ability to perform addition and subtraction. This suggests that sequence continuation, addition, and subtraction tasks share circuits. As shown in Table 3, we find that ablating the sequence continuation component heads destroy the ability to perform double-digit arithmetic for 92-100% of our experiments, while randomly selected ablated heads only do so approximately 20-33% of the time.
Table 3: Percentage Destroyed after Ablation. Percentage Destroyed denotes the percentage of prompts with incorrect completion after ablation of the component sets of that column; the higher the score, the more effective the ablation. Each prompt type uses 50 prompts. Results for more prompt types are given in Table 6.
@@
8 Conclusion
9 Limitations
Methodology Assumptions. As mechanistic interpretability is an early research field, there may be methodological assumptions that these findings depend on which may be subject to change as the research field progresses.
Dataset Size. As there are only twelve months, the number of possible continuing sequences was limited. Additionally, even if months were not used, GPT-2 also has limited prediction ability for number word sequences, as detailed in Appendix B. However, our dataset size for the months continuation task fully captures all of the months; in contrast, a small dataset size brings more issues when it does not capture all of the true distribution.
Future Work. Future work can build upon this paper by tackling open questions related to feature-level analysis of sequence continuation tasks, and by examining components exclusive to certain tasks to see if they handle mapping between abstract representations to specific tasks. Another research problem involves analyzing the effects of model editing on shared, entangled circuits. Approaches to this include quantifying the relationship between circuit entanglement and editing impact which may be done via embedding space projection (Dar et al., 2022), modifying a sub-circuit used for a sub-task $S$ and observing if the ability to recognize $S$ in multiple tasks is destroyed, and utilizing methods such as model steering to edit a task $S$ to perform a similar task $S^{\prime}$ Turner et al. (2023); Merullo et al. (2023).
Acknowledgements
We are grateful to Torr Vision Group (TVG) and Apart Lab members for feedback on the previous version, specially, Jishnu Mukhoti, Luke Marks, Michelle Lo and Ashkan Khakzar for comments and feedback and Clement Neo for assistance with Figure 1.
Appendix A Computational Resources and Packages
@@
The code for the experiments was written in Python, utilizing the TransformerLens package, and were run on an A100.
Appendix B GPT-2 Experimental Setup Details
@@
Appendix C Llama-2 Experimental Setup Details
@@
Dataset. For Llama-2, we ran iterative node pruning on samples consisting only of sequence members, and ran attention patterns on samples consisting of both sequence and non-sequence members. This is because for Llama-2, the main focus was on finding similar components between the sequence continuation sub-circuits of GPT-2 and Llama-2, rather than finding more minimal circuits, which would be beyond the computational resources we allocated. Additionally, the numbers 10 to 12 are represented as two tokens (eg. ’10’ is ’1’ and ’0’) in Llama-2-7b; as we evaluate only on prompts of the same length and use single token answers, we only use numbers 1 to 9 in the Numerals dataset.
@@
Iterative Methods. Given that Llama-2 is larger than GPT-2, we only use on backwards sweep, which we found to be enough for finding the most important components. We also try modifying the iterative component ablation method to first ablate all the heads of an attention layer to see if this makes the model performance perform below the threshold; if it does, we remove all the heads of that attention layer from the candidate circuit. While this method variation is faster if many attention layers are unnecessary, it has the downside in which removing certain unimportant heads may increase performance, which may, in summation, offset the performance degrading abilities of important heads. This means that there may be some important heads which may be missed by this method. However, we found that the most important attention heads were not missed by this method, and the end results were often comparable to the original version of our iterative methods. The results in this paper do not use this modification.
Appendix D Iterative Path Patching for Edges
Path patching is a different type of patching that allows for a more precise analysis of an intervention’s effect on a particular path Goldowsky-Dill et al. (2023). It can be performed by ablating component interactions, measuring the effect of one component on another.
After finding circuit nodes, we utilize path patching to obtain interactions (edges) between them. Edges denote nodes with high effects on other nodes  [^9] As the residual stream allows for indirect effects, edges may be between components at non-adjacent layers,. We apply a form of iterative path patching which works backwards from the last layers by finding earlier components that affect them. First, we ablate the nodes pruned from iterative node pruning. Then, we ablate one candidate edge of the unablated nodes at a time. Using the same order as the backward sweep, we take a node as a receiver and find the sender nodes that have an important effect on it. If patching the effect of sender A on receiver B causes the model performance to fall below threshold $T_{e}$, the edge is kept; else, it is removed.
Appendix E Functionality Method Details
Component Output Scores Details. To continue from Section §3, we employ the heads’ output projection ($\mathbf{OV}$) matrices to examine the attention head outputs written to the residual stream by the OV circuit. For example, we can check if a head is copying tokens, a behavior introduced as copy scores by Wang et al. (2022). Copy scores measure how well a head reproduces a token from the input. Similar to Gould et al. (2023), who analyzed successor heads in greater detail than our paper, we modify this method to obtain the component output score, or OV score, which follows a similar principle but measures how many prompts have a keyword token are in the output. To calculate these scores, we multiply the state of the residual stream after the first MLP layer at the last token with the OV matrix of the attention head of interest. This result is unembedded and layer normalized to get logits. If the keyword is in the top-5 of these logits, +1 is added to the score. Finally, we divide the total score by the total number of keywords across all prompts to obtain a percentage which we call the OV score. Specifically, we use the keyword as the integer $I+1$, given integer $I$ as the last token in a sequence, to obtain an OV score known as the successor score. In this paper, we use all sequence members of the prompt as keywords.
Appendix F GPT-2 Small Connectivity Details
As seen in both Figure 8 and in Table 8 in Appendix K, in which only head 5.0 of the Numerals circuit is not part of the Number Words circuit, the Numerals circuit is nearly a subset of the Number Words circuit. This suggests that the Number Words circuit uses the Numerals circuit as a sub-circuit, but requires additional components to make accurate predictions.
Numerals Task NumWords Task Months Task
Numerals Circuit 81.01% 48.41% 113.52%
Number Words Circuit 87.35% 81.11% 103.64%
Months Circuit 43.74% 32.36% 80.30%
IOI Circuit -6.70% -15.82% -9.20%
Appendix G GPT-2 Small Functionality Details
G.1 Last Token Sequence Detection Head
G.2 Successor Head Scores
To check that head 9.1 outputs next sequence tokens, we study its component output scores. Table 5 shows that given a numeral token I as input (eg. 1), head 9.1 often outputs a token I+1 or higher (eg. 2). For numerals between 1 and 97, its successor score is 87.37%, while its copy score is 59%. We also note that the successor scores of most heads are low, with an average of 3.29%, and that head 9.1 has the highest successor score. Thus, it seems to function as a "successor head".
This is reinforced by the successor score for number words, which is 90.63%, while the average for all attention heads is 2.97%. Although head 9.1 does not appear to output months given any month token, we observe something peculiar: 9.1 is the only head that will output the next rank given a month (eg. given "February", output "third", and its "next rank given month" score is 31.25%. This appears to be related to how months can be mapped onto ranks.
Token Top-3 Tokens after Unembed
‘78’ ’ 79’, ‘80’, ‘81’
‘six’ ’ seventh’, ’ eighth’, ’ seven’
‘August’ ‘ighth’, ‘eighth’, ‘ninth’
Appendix H Llama-2-7B Functionality Details
Other Attention Patterns. Figure 7 shows attention patterns attention 16.0 It may a slight resemblence to the "last token sequence detection head" 7.11 in GPT-2 Small, but it is not clear from this plot what the role of this head is.
Figure 7: Llama-2-7B attention pattern for Head 16.0. Tokens appears to attend to the last period of the first instance of the pattern (<object> lost in Month.) However, it is not clear from this plot what the role of this head is.
Appendix I Ablation on Math-Related Prompts
Experimental Setup. We run ablation experiments on various math-related prompt types only with Llama-2-7B as GPT-2 Small is limited in its ability to correctly complete many math-related prompt types. Zero ablation is performed for these experiments, in which the components being knocked out are replaced with an activation of 0, rather than mean activations from a corrupted dataset. This is because for certain prompts like for natural language, it is not always clear what type of corruption to choose, so zero ablation is easier to work with.
For random ablation on intervaled sequences and arithmetic, a sequence continuation component set for a task is around 80-90 attention heads out 1024 total attention heads. Thus, we choose to ablate 100 attention heads. For the randomly selected component sets for all prompt types, we do not select attention heads that belong to the the intersection set of attention heads from the 3 circuit found for the three sequence continuation tasks. In other words, the Numerals circuit component set has 86 attention heads, and taking the intersection of this component set with the component sets from the Number Words and Months circuits yields an intersection set of 16 attention heads. None of these 16 heads are not selected when obtaining attention heads for the randomly selected component sets.
Overall, we note that these tasks are not solely dependent on the sequence continuation component sets we find, which means random ablations may choose components that are crucial for the task. For instance, the random ablation score for the arithmetic may likely be lower, as some of its runs may be ablating heads that are involved in arithmetic.
Math-Related Word Problems. The following experiments provide some evidence that the sequence continuation circuits are not specific to just sequence continuation, but can extend to prompts involving sequence continuation reasoning. More experiments on larger datasets and more prompts can determine if statement holds in more cases. In summary, we evaluate on the following prompt types, with type (2) and (3) being prompt templates:
  1. 1.
    "What are the months in a year? Give all of them as a list. Be concise."
  2. 2.
    "If this month is X, and Y months pass, what month is it? Answer: "
  3. 3.
    "If today is the Xth of (month M), what date will it be in Y days? "
(1) "What are the months in a year? " We observe the effects of ablating sequence continuation task circuits on the prompt, "What are the months in a year? Give all of them as a list. Be concise." To evaluate the model’s answer, we allow the model to generate an output of 50 tokens. We apply the iterative pruning algorithm to obtain component sets for the three tasks. When ablating any of the 3 sequence continuation component sets, which contain around 80 heads each (eg. the Numerals circuit has 86 heads), the model loses the ability to correctly complete the answer. In particular, ablating the Numerals attention head set will make the model stop the list at June, ablating the Number words attention head set will make the model stop the list at February, and ablating the Months attention head set will make the model state the output as "1. January 2. February 2. February 3. March 3. April 3. May 3. August 3. August", listing each item on a new line. We note that the Numerals and Number Words head sets do not particularly show strong evidence for sequence destruction, but the Months head set clearly shows the model loses the ability to correctly complete the sequence. In contrast, choosing 86 random heads that do not overlap with the Numerals circuit component set (which has 86 heads) will allow the model to retain the correct ordering for 47 out of 50 random samples of 86 random heads. We evaluate these outputs using GPT-4.
(2) "If this month is X, and Y months pass, what month is it? Answer: " We study ablation on prompts of the format, "If this month is X, and Y months pass, what month is it? Answer: ". We evaluate on 43 prompts generated using this template that the model obtains the correct answer on when unablated.
We find that ablating the 3 functionally important heads yields a percentage destroyed score of 37.21%, while using 20 randomly sampled component sets of three attention heads yields a mean destruction score of 12.62%. For each run, we calculate the percentage destroyed score, and take the mean over every run. We also find that the three functionally important attention heads 5.25, 16.0 and 20.17 may work together, because if one of them is not present, the model still retains its correct completion ability.
To evaluate if the model obtains the correct answer, we notice that for each prompt, the model may answer in different formats. For instance, for the prompt "If this month is December, and 12 months pass, what month is it? Answer: ", the model answers: "12 months have passed, so it is now December again". However, for the prompt, "If this month is April, and 9 months pass, what month is it? Answer: ", the model answers: "If it is April and 9 months pass, then it will be January". Thus, we determine if the model obtains the correct answer by extracting all the months it answers with for each prompt. Then, we manually inspect both the full answer and the extracted month to determine if the extracted month is actually consistent with the model logically stating the extracted month as the answer, or if the model is simply "mentioning" that month. We find that for the prompts we use, the model is actually logically stating the extracted month as the answer. We allow the model to generate 15 tokens in its output. We found that for most prompts, this is enough for the model to correctly obtain the right answer, as shown for a sample in Table 7.
We also found that giving an instruction such as "Be concise" or a pattern we expect the model to complete in, such as "If this month is March, and 2 months pass, what month is it? Answer: May. ", often makes the model give wrong answers. Instead, we found that just letting the model generate from the question itself would yield often the correct answer.
(3) "If today is the Xth of (month M), what date will it be in Y days? " The correct answer for the majority of prompts is given in "Month Day" format (eg. January 10th). Thus, we the model to generate 5 tokens in its output.
We use 41 prompts that the model completes correctly, and discover that we only need to ablate the 3 functionally important heads for sequence continuation- 5.25, 16.0, and 20.17- to disrupt these word prompt correct completion. In contrast, ablating just 3 heads at random often does not destroy the model’s correct completion abilities. In detail, we find that ablating our three important heads destroys the correct completions for 24.4% of prompts, while randomly ablating on 3 heads, on average over 10 randomly sampled component sets, destroys 9.2% of prompts. We also notice that ablating the 3 functionally important heads often shifts the day by just a few days, rather than changing the month.
Overall Analysis: We note that the last two prompt types are still sensitive to ablating any 3 heads of the model, suggesting that the model may require most of the heads of the model to work properly for these tasks, or that its predictions for the right answer are not as strong (having more weight on the right answer over other tokens). The latter is likely true given that the model does not always obtain the correct answer for these prompt templates; given 100 prompts, we selected 43 and 41 prompts, respectively, for the previous two templates as these were the ones the model correctly completed on. The destruction score for the 3 heads is also closer to the random ablation score. Thus, these prompt types are relatively weaker evidence for the effect of the discovered sequence continuation circuits on math-reasoning word problems; nonetheless, there is some noticeable effect, so we have reported its results in this Appendix. Using a better (fine-tuned or larger) model that has improved reasoning skills can mitigate this issue. More randomly selected samples may also be used.
In contrast to the math-related reasoning prompts discussed above, we find that a "simpler" sequence continuation task such as "1 2 3" requires more heads- around 40 to 80- for ability destruction to occur for most prompts. This may suggest that more complex prompts are more sensitive to requiring most components be healthy, while certain simpler tasks do not require this and are only destroyed when many components are destroying, possibly suggesting that these large circuit component sets we discovered may contain many backup sub-circuits within them.
Prompt Type Numerals NumWords Months Random 100
Single-Digit Addition 62% 46% 92% 15.6%
Single-Digit Subtraction 72% 46% 74% 12.4%
Spanish Words Sequence Continuation. We looked into Llama-2-7B’s capabilities for sequence continuation across various languages, but found it to be limited. For instance, the model is able to complete "uno dos tres" correctly, but will output numeral "6" for prompt "dos tres cuatro cinco" instead of "seis". Thus, we only study two prompts for Spanish number word continuation that Llama-2-7B successfully completes on: "uno dos tres" and "seis siete ocho", and two prompts for Spanish months continuation: "enero, febrero, marzo". Lastly, we find the model can also correctly complete Spanish months continuation prompts when asked in natural language question formats, and find that ablating the sequence continuation circuits also affects these inputs.
Appendix J All Three Circuits Diagram for GPT-2 Small